OpenStudy (anonymous):

rewrite with only sin x and cos x. cos 3x

OpenStudy (anonymous):

you can rewrite it: \(\cos(3x) = \cos(2x+x)\)

OpenStudy (anonymous):

then you use sum formula

OpenStudy (anonymous):

yeah, i got it to cos\[\cos 2x(\cos x- \sin 2x \sin x)\]

OpenStudy (anonymous):

now do same with cos 2x and sin 2x: \(\cos (2x) = \cos(x+x)\)